### v2

# Defending Against Alignment-Breaking Attacks via Robustly Aligned LLM

Abstract
Recently, Large Language Models (LLMs) have made significant advancements and are now widely used across various domains. Unfortunately, there has been a rising concern that LLMs can be misused to generate harmful or malicious content. Though a line of research has focused on aligning LLMs with human values and preventing them from producing inappropriate content, such alignments are usually vulnerable and can be bypassed by alignment-breaking attacks via adversarially optimized or handcrafted jailbreaking prompts. In this work, we introduce a Robustly Aligned LLM (RA-LLM) to defend against potential alignment-breaking attacks. RA-LLM can be directly constructed upon an existing aligned LLM with a robust alignment checking function, without requiring any expensive retraining or fine-tuning process of the original LLM. Furthermore, we also provide a theoretical analysis for RA-LLM to verify its effectiveness in defending against alignment-breaking attacks. Through real-world experiments on open-source large language models, we demonstrate that RA-LLM can successfully defend against both state-of-the-art adversarial prompts and popular handcrafted jailbreaking prompts by reducing their attack success rates from nearly 100% to around 10% or less.
WARNING: This paper contains unsafe model responses. Reader discretion is advised.
1 INTRODUCTION
Trained on a wide range of text data from the internet, Large Language Models (LLMs) have exhibited exciting improvement in their generalization capabilities (OpenAI, 2023; Touvron et al., 2023b) and widespread application in various domains such as finance (Wu et al., 2023), law (Nguyen, 2023), and healthcare industry (Thirunavukarasu et al., 2023). While LLMs have showcased impressive potential, a rising concern is that they can also be maliciously utilized to generate content deviating from human values (e.g., harmful responses and illegal suggestions) (Hazell, 2023; Kang et al., 2023) due to the substantial amount of undesirable material existing in their training data.
Figure 1: An illustration of alignment-breaking attack: an aligned LLM gives unsafe responses to malicious requests with adversarial prompts.
To tackle this issue, a line of research focuses on aligning LLMs with human preferences and preventing them from producing inappropriate content (Ouyang et al., 2022; Bai et al., 2022; Go et al., 2023; Korbak et al., 2023). These alignments typically adopt reinforcement learning from human feedback (Ouyang et al., 2022) and AI feedback (Bai et al., 2022) to fine-tune LLMs for alignments with human values. Despite these efforts, an emerging class of jailbreak attacks can still bypass the alignment and elicit harmful responses from LLMs (Yuan et al., 2023; Shen et al., 2023; Wei et al., 2023; Zou et al., 2023). These alignment-breaking attacks manually craft adversarial prompts by designing elaborate role-play (Shen et al., 2023) or simply asking the LLM to give the response starting with “Absolutely! Here’s” (Wei et al., 2022). Moreover, automatic jailbreak prompt generation methods have also been developed through dialogue encryption (Yuan et al., 2023) or the combination of greedy and gradient-based search methods (Zou et al., 2023). Figure 1 shows an example that a malicious question appended with an adversarial prompt could successfully break the safety alignment. Recently, (Zou et al., 2023) have demonstrated that jailbreak attempts could be highly effective and transferable across different LLMs. This phenomenon suggests that existing safety alignment is far from robust to defend against carefully crafted adversarial prompts.
Till now, few attempts have been made to design dedicated mechanisms for defending alignment-breaking attacks. A rudimentary defense currently employed relies on external tools to re-assess the potential harm of the LLM responses. For instance, it could feed every potential response from the target LLM into a third-party LLM to determine whether the response is harmful or not (Helbling et al., 2023). While this strategy enables filtering out possible harmful responses, there are several major drawbacks limiting its practicability: 1) Existing LLMs are very sensitive to harmful keywords appeared in the input, and have a high propensity to misclassify benign content as harmful, even when the entire sentence is not talking about any harmful behavior (e.g., stating news or providing guidance/warnings). This could lead to a high false-positive rate in harmful content detection; 2) The method heavily relies on the performance of the LLM used as a harmful discriminator, while the LLM itself is not designed to be an accurate harmful discriminator. The basis for its decisions remains ambiguous, implying that the harmful evaluation process could be opaque; 3) There are more types of alignment that can not be simply summarised as “harmful” (e.g., privacy, ethics, human values etc), thus this type of approach cannot cover such cases simultaneously. Given the wide range of applications where LLMs could be utilized, finding an effective and practical defense against potential alignment-breaking attacks is both urgent and challenging.
In this work, we design a Robustly Aligned LLM (RA-LLM) to defend against potential alignment-breaking attacks, which is built upon an already aligned LLM and makes the existing alignments less prone to be circumvented by adversarial prompts. Specifically, our key idea is that although an aligned LLM can, to some extent, identify if the input request is benign or not, we cannot directly rely on that as it may not be robust. We consider an input request to be benign, only if we randomly drop a certain portion of the request and the LLM still thinks it is benign in most cases. Intuitively, such a random dropping operation would invalidate the adversarial prompts in alignment-breaking attacks, which are usually sensitive to small perturbations; on the other hand, the chances for the LLM to reject benign requests are relatively low, even after random dropping. Therefore, such a mechanism naturally leads to a robustly aligned LLM.
Note that our RA-LLM does not require any external “harmful” detectors, instead, our strategy only relies on the existing alignment capability inside the LLM. Due to the same reason, our approach is not limited to any specific type of alignment (e.g., harmful), but robustifies all existing model alignments. Furthermore, we provide a theoretical analysis to verify the effectiveness of our proposed RA-LLM. Our experimental results on open-source large language models demonstrate that RA-LLM can successfully defend against both state-of-the-art adversarial prompts and popular handcrafted jailbreaking prompts by reducing their attack success rates from nearly 100% to around 10% or less.
2 RELATED WORKS
Aligning LLMs with Human Preferences
Foundational large language models are pre-trained on extensive textual corpora (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a), which equips LLMs with world knowledge and facilitates their deployment in professional applications. Despite their excellent performance, LLMs suffer from generating outputs that deviate from human expectations (e.g., harmful responses and illegal suggestions) due to the significant amount of inappropriate content existing in unfiltered training data. To tackle this issue, a line of work focuses on aligning LLMs with human values (Xu et al., 2020b; Ouyang et al., 2022; Bai et al., 2022; Go et al., 2023; Korbak et al., 2023). Specifically, Ouyang et al. (2022) align the LLM by using reinforcement learning from human feedback (RLHF (Christiano et al., 2017; Stiennon et al., 2020)) to fine-tune pre-trained LLM with human preferences as the reward signal, which reduces the generation of toxic content. Bai et al. (2022) train a less harmful system through the specification of a short list of principles and further improve the human-judged performance by introducing chain-of-thought style reasoning (Wei et al., 2022) in both supervised-learning and reinforcement-learning stage. Go et al. (2023) consider aligning LLMs as approximating a target distribution representing some desired behavior and accordingly propose a new framework for fine-tuning LLMs to approximate any target distribution through f-divergences minimization. In addition to aligning LLMs in the fine-tuning stage, Korbak et al. (2023) propose pertaining LLMs with alternative objectives that guide them to generate text aligned with human preferences and significantly reduce the rate of generating undesirable content by using conditional training (Keskar et al., 2019).
Alignment-breaking Attacks and defenses in LLMs
Although various alignment strategies have been developed to steer LLMs to generate content complying with human ethical principles, an emerging class of alignment-breaking attacks (i.e., jailbreak attacks) can still bypass safeguards and elicit LLMs to generate harmful and toxic responses (Wolf et al., 2023; Li et al., 2023; Shen et al., 2023; Yuan et al., 2023; Wei et al., 2023; Zou et al., 2023), which poses significant threats to the practical deployment of LLMs. In particular, inspired by traditional computer security, Kang et al. (2023) adapt obfuscation, code injection/payload splitting, and visualization attacks to the LLMs, leading to the generation of content containing hate speech, phishing attacks, and scams. Wei et al. (2023) hypothesize that competing objectives and mismatched generalization are two failure modes of safety training in LLMs and craft effective jailbreak attacks by leveraging the two failure modes. Instead of manually crafting adversarial prompts, Zou et al. (2023) automatically produce transferable adversarial suffixes by using greedy and gradient-based search methods to maximize the probability of generating an affirmative response. Yuan et al. (2023) bypass the safety alignment through dialogue encryption. Shen et al. (2023) systematically analyzes the characteristics of jailbreak prompts in the wild and presents that jailbreak prompts have evolved to be more stealthy and effective with reduced length, increased toxicity, and semantic shift. Note that some concurrent works also aim to defend against alignment-breaking attacks: Kumar et al. (2023) provides a verifiable safety guarantee by enumerating all possible partially erased input and using a safety filter to identify the harmfulness of the input content. Jain et al. (2023) propose to detect adversarial prompts by checking if the perplexity of the prompt is greater than a threshold.
Traditional Text Adversarial Attack and Defenses
Traditional text adversarial attacks primarily focus on text classification tasks and aim to force target models to maximize their prediction error by adversarially perturbing original text (Ebrahimi et al., 2017; Jin et al., 2020; Li et al., 2018; Maheshwary et al., 2021; Ye et al., 2023). The adversarial perturbation could be crafted by performing character-level transformation (Gao et al., 2018) or replacing original words with their synonyms while maintaining semantics and syntax similar (Alzantot et al., 2018). The generation of adversarial examples could be categorized into the “white-box” setting and the “black-box” setting according to the extent of access to the target model (Xu et al., 2020a). As a representative white-box method, HotFlip  (Ebrahimi et al., 2017) uses the gradient information of discrete text structure at its one-hot representation to construct adversarial examples. In the black-box setting, Li et al. (2018); Jin et al. (2020); Ren et al. (2019) leverage the prediction score distribution on all categories to craft adversarial text without the guidance of parameter gradients. Maheshwary et al. (2021) focus on a more realistic scenario where attackers only know the top-$1$ prediction and propose using population-based optimization to construct adversarial text. Ye et al. (2022) follow the same scenario and employ the word embedding space to guide the generation of adversarial examples.
To defend against adversarial attacks, a body of empirical defense methods has been proposed. In particular, adversarial-training-based methods (Miyato et al., 2016; Zhu et al., 2019) incorporate adversarial perturbations to word embeddings and robustly train the model by minimizing the adversarial loss. Zhou et al. (2021); Dong et al. (2021) utilize adversarial data augmentation by replacing the original word with its synonyms to make the model robust to similar adversarial perturbations. These methods gain empirical success against adversarial attacks. To provide provable robustness against adversarial word substitutions, Jia et al. (2019) use certifiably robust training by training the model to optimize Interval Bound Propagation (IBP) upper bound. Shi et al. (2020) adopt linear-relaxation-based perturbation analysis (Xu et al., 2020c) to develop a robustness verification method for transformers. Zeng et al. (2023) propose a certifiably robust defense method based on randomized smoothing techniques (Cohen et al., 2019).
3 Our Proposed Method
In this section, we introduce the proposed Robustly Aligned LLM for defending alignment-breaking attacks. Before heading into details, we first discuss the threat model that is focused on in this paper.
3.1 Threat Model
An alignment-breaking attack seeks to bypass the security checks of an aligned LLM by introducing adversarial prompts adhered to an original malicious question. Let ${\mathbf{x}}$ denote a malicious question and ${\mathbf{p}}_{\text{adv}}$ represent the adversarial prompt generated by the alignment-breaking attack. Let ${\mathbf{x}}_{\text{adv}}={\mathbf{x}}\oplus{\mathbf{p}}_{\text{adv}}$ denote the entire input (malicious question and the adversarial prompt) where $\oplus$ denotes the insertion operation. While most existing attacks typically place the adversarial prompts at the end of the request Zou et al. (2023), we actually consider a more general case where the adversarial prompt could also be inserted in front of the malicious question or be integrated in the middle.
We also assume that the target LLM $f(\cdot)$ is an already aligned LLM that has a certain ability to reject commonly seen malicious requests. In other words, when the malicious question ${\mathbf{x}}$ is directly input into the target LLM $f(\cdot)$, it will, in most cases, deny answering such a question by outputting a response similar to “I am sorry, but I cannot talk about [a malicious request]…”. On the contrary, the alignment-breaking attacker’s goal is to break the existing alignment of the target LLM by finding an adversarial prompt ${\mathbf{p}}_{\text{adv}}$, so that ${\mathbf{x}}_{\text{adv}}={\mathbf{x}}\oplus{\mathbf{p}}_{\text{adv}}$ will mislead the LLM to provide an affirmative answer Zou et al. (2023) to such a malicious question, e.g., “Sure, here is how to do [a malicious request]…”.
3.2 Our Proposed Method
Our motivation builds upon the fact that the target LLM has already been aligned and is able to reject commonly seen malicious requests. To be more specific, we can build an alignment check function $\text{AC}(\cdot)$ based on the aligned LLM $f(\cdot)$: return Fail when detecting typical aligned text in the output of $f(\cdot)$ such as “I am sorry, but I cannot answer this …”, and return Pass otherwise [^1] This definition of $\text{AC}(\cdot)$ is quite vague but we will provide more details on how to implement it in practice in Section 3.3.. Given the alignment check function $\text{AC}(\cdot)$, one can then construct a “hypothetical” LLM by
$f^{\prime}({\mathbf{x}})=\left\{\begin{aligned} &\text{Reject the response},\text{ if }\text{AC}(f({\mathbf{x}}))=\text{Fail}\\ &f({\mathbf{x}})\ \ \ \ \ \ \ \qquad\qquad,\text{ if }\text{AC}(f({\mathbf{x}}))=\text{Pass}\end{aligned}\right.$ (1)
where $f^{\prime}({\mathbf{x}})$ denotes the “hypothetical” LLM constructed by using the alignment check function $\text{AC}(\cdot)$. While $f^{\prime}({\mathbf{x}})$ seems “useless” as it gives the same result as $f({\mathbf{x}})$ in practice, it showcases how one can construct a new aligned LLM using an alignment check function.
Robust Alignment Check Function
One thing to notice here is that the previously defined alignment check function $\text{AC}(\cdot)$ only relies on the existing alignments inside on the target LLM. However, the existence of alignment-breaking attacks such as the adversarial prompts Zou et al. (2023) has proved that such alignment checking is not robust: it can be easily manipulated and circumvented by carefully designed perturbations or suffix prompts. Therefore, it is natural to think about how we can design a robust alignment check function that could strengthen the alignment check capabilities of an aligned LLM, without finetuning or modifying the model itself.
Our intuition here is very straightforward: since the existing alignment check function $\text{AC}(\cdot)$ is not robust enough for alignment-breaking prompts, the Pass decision directly returned by it cannot fully convince us that the request is benign, instead, we need a stronger evidence and a stricter check. Specifically, we consider an input request to be benign, only if we randomly drop a certain portion of the request and the corresponding response can still pass the alignment check function $\text{AC}(\cdot)$ in most cases. To translate this requirement into mathematical formulations, we define the following Robust Alignment Check function $\text{RAC}(\cdot)$ based on the aligned LLM $f(\cdot)$ and the alignment check function $\text{AC}(\cdot)$ :
$\text{RAC}({\mathbf{x}})=\left\{\begin{aligned} &\text{Fail},\ \text{ if}\ \text{AC}(f({\mathbf{x}}))=\text{Fail}\\ &\text{Fail},\ \text{ if}\ \mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U(p)}(\text{AC}(f([{\mathbf{x}}]_{{\mathbf{r}}}))=\text{Fail})>t\\ &\text{Pass},\ \text{ otherwise}\end{aligned}\right.$ (2)
where ${\mathbf{r}}$ refers to the uniformly sampled indices mask to indicate kept tokens, $U(p)$ refers to the distribution of possible masks after uniformly dropping $p$ portion of the indices (without replacement), and $[{\mathbf{x}}]_{{\mathbf{r}}}$ denotes the kept indices ${\mathbf{r}}$ inside ${\mathbf{x}}$ after the dropping operation. Essentially, for an input ${\mathbf{x}}$ with length $L$, every possible $[{\mathbf{x}}]_{{\mathbf{r}}}$ only contains $(1-p)L$ tokens indexed by ${\mathbf{r}}$.
Eq. 2 states that the robust alignment check function $\text{RAC}(\cdot)$ not only requires the original response $f({\mathbf{x}})$ to show no sign of being aligned but also requires the response after random dropping still shows no sign of being aligned in most cases. On the contrary, if $\text{AC}({\mathbf{x}})$ already fails or over a certain ratio (e.g., $>t$) of responses from the randomly dropped input fails to pass AC, $\text{RAC}(\cdot)$ will also fail it. Therefore, it is easy to see that such a design certainly helps us build a more robust alignment check function compared to $\text{AC}(\cdot)$.
Based on the robust alignment check function $\text{RAC}(\cdot)$, we can further construct a robustly aligned LLM by simply replacing the vanilla alignment check function $\text{AC}(\cdot)$ with $\text{RAC}(\cdot)$ in Eq. (1):
$f_{\text{rob}}({\mathbf{x}})=\left\{\begin{aligned} &\text{Reject the response},\text{ if }\text{RAC}(f({\mathbf{x}}))=\text{Fail}\\ &f({\mathbf{x}})\ \ \ \ \ \ \ \qquad\qquad,\text{ if }\text{RAC}(f({\mathbf{x}}))=\text{Pass}\end{aligned}\right.$ (3)
By this simple reconstruction of alignment check function, we can build a robustly aligned LLM without necessitating extra resources or retraining of the entire model. Figure 2 illustrates the effect of our proposed RAC when facing malicious or benign requests.
Figure 2: An illustration of our RA-LLM when facing malicious requests with adversarial prompts (Left) and benign requests (Right).
3.3 Practical Designs
Algorithm 1 Robustly Aligned LLM
Input: aligned LLM $f$, alignment check function AC, original input ${\mathbf{x}}$.
1:  if $\text{AC}(f({\mathbf{x}}))=\text{Fail}$  then
2:     Reject the request
3:  else
4:     for $i=1,2,\cdots,n$ do
5:        Randomly sample a mask ${\mathbf{r}}_{i}\sim U(p)$
6:        $s_{i}=\mathbbm{1}\{\text{AC}(f([{\mathbf{x}}]_{{\mathbf{r}}_{i}}))=\text{Fail}\}$
7:     end for
8:     if $(1/n)\sum_{i=1}^{n}s_{i}>t$ then
9:        Reject the request
10:     else
11:        Return $f({\mathbf{x}})$
12:     end if
13:  end if
Now let’s delve into the practical designs of our proposed robustly aligned LLM, which essentially approximates $f_{\text{rob}}(\cdot)$ mentioned above. The detailed steps of the constructed robustly aligned LLM are summarized in Algorithm 1.
Approximation of $\text{AC}(\cdot)$
Previously, we vaguely defined the alignment check function $\text{AC}(\cdot)$ as returning Fail when detecting typical aligned output while returning Pass otherwise. In practice, we approximate this alignment check function through prefix checking: we observed that various aligned outputs often share similar prefixes such as “I can not”, “I’m sorry”. Therefore, we can build a prefix set and if any prefix in the set appears in LLM’s output, the alignment check function $\text{AC}(\cdot)$ returns Fail; otherwise, it returns Pass. Note that we are only inspecting the prefix. For this purpose, we only need to generate a certain number of tokens (e.g., 10) for robust alignment checking. This could largely reduce our computational overhead [^2] Further discussion on computational costs can be found in Section 4.5..
Monte Carlo Sampling
It is practically infeasible to obtain the exact value for the probability of $\mathop{{\mathbb{P}}}_{{\mathbf{r}}\sim U(p)}(\text{AC}(f([{\mathbf{x}}]_{{\mathbf{r}}}))=\text{Fail})$, as it would require us to enumerate all possible random dropping cases and is computationally intractable. Therefore in practice, we conduct Monte Carlo sampling to approximate the true probability: we randomly sample $n$ indices masks to obtain $n$ versions of the input request with random dropping; we then solicit the LLM’s responses for these $n$ requests, and count the frequency of cases when the alignment check function $\text{AC}(\cdot)$ gives Fail decisions.
The Practical Choice of $t$
Another important choice is the threshold $t$ used in practice. In particular, one seemingly logical choice is setting $t\to 0$ such that whenever $\text{AC}(\cdot)$ detects any failure case from the randomly dropped request, $\text{RAC}(\cdot)$ directly fails the request. However in practice, such a setting could be too extreme as the randomness introduced in the dropping operations might also affect the LLM response on benign inputs: random dropping might occasionally lead to the loss of essential information, and under such circumstances the LLM might also generate responses similar to the typical alignment responses. For example, “Do you like apples?” could become “Do you apples?” after random dropping, leading the LLM to express an inability for answering this unclear question. This could potentially be mis-detected as Fail by $\text{AC}(\cdot)$, and if the threshold $t\to 0$, it will lead to Fail by $\text{RAC}(\cdot)$ and be rejected by our robustly aligned LLM. Therefore, in practice, instead of setting the threshold $t$ as zero, we keep a relatively small threshold.
3.4 Theoretical Analysis
In this section, we theoretically analyze the proposed robustly aligned LLM and see when it provides a more robust alignment compared to the original LLM when facing alignment-breaking attacks.
Our theorem is based on the analysis on the robust alignment check function RAC. We will show that RAC is more robust for the aligned malicious text $x$ with any adversarial prompt ${\mathbf{p}}_{\text{adv}}$ of length $M$ and it can be inserted into any position (e.g., in front, back, or middle of ${\mathbf{x}}$).
Theorem 3.1.
[…]
The proof of Theorem 3.1 is provided in Appendix A. Theorem 3.1 provides an analysis on when our robustly aligned LLM could reject the request from an alignment-breaking attack while the original LLM actually fails to defend against such adversarial prompts. Specifically, given a particular malicious input ${\mathbf{x}}$ whose response has been aligned by the target LLM $f(\cdot)$, although it is impossible for us to know what kind of adversarial prompt the attacker would use, or which position the attacker would insert the adversarial prompt to, as long as we have $\mathop{\min}\limits_{j}\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U(p)}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})>t+c$, then any alignment-break attack ${\mathbf{x}}_{\text{adv}}$ composed by ${\mathbf{x}}\oplus{\mathbf{p}}_{\text{adv}}$ will be rejected by our robustly aligned LLM.
4 Experiments
In this section, we aim to validate the efficacy of our RA-LLM from two aspects: 1) RA-LLM can effectively reduce the attack success rate of adversarial prompts; 2) RA-LLM minimally affects the outputs of benign samples. In the following, we first introduce our experimental settings and give a detailed analysis of our experimental results and ablation study.
4.1 Experimental Settings
Dataset
We evaluated our approach on two datasets: AdvBench (Zou et al., 2023) and MS MARCO dataset (Nguyen et al., 2016). The AdvBench dataset contains two types of data, corresponding to Harmful Strings Attack and Harmful Behaviors Attack respectively. Specifically, The data used for Harmful Strings Attack consist of 500 strings related to harmful or toxic content, such as threats, discriminatory remarks, methods of crime, and dangerous suggestions, etc. The data used for Harmful Behaviors Attack consists of 500 questions that can entice the LLM to produce harmful outputs, with topics similar to Harmful Strings. MS MARCO is a question-answering dataset, where all questions originate from real user queries on Bing. We sampled 150 pieces of data from each of these three datasets for our experimental evaluation.
Attack Setting
We mainly evaluate our defense under the state-of-the-art alignment-breaking attack: the Harmful Behaviors Attack proposed by (Zou et al., 2023). The goal of Harmful Behaviors Attack is to induce the LLM to respond effectively to malicious queries, which normally should be rejected by the aligned LLMs. Harmful Behaviors Attack aims to bypass the protective measures of aligned LLMs and entice them to generate harmful content. We calculated all adversarial prompts using the default hyperparameters provided in (Zou et al., 2023).
4.2 Experimental Results
In Table 1, we present the experimental results on two attack modes of the Harmful Behaviors Attack: Individual Attack and Transfer Attack, on Vicuna-7B-v1.3-HF and Guanaco-7B-HF models. Individual Attack aims to directly optimize adversarial prompts for specific models and specific malicious requests, while Transfer Attack aims to optimize generic adversarial prompts across multiple models and malicious requests. We tested both the original aligned LLM and our robust aligned LLM using benign requests and malicious requests with adversarial prompts. Subsequently, we evaluated whether these inputs activated the alignment mechanism based on the output of the LLM.
Table 1: The benign answering rate and attack success rate of the original LLM and our robustly aligned LLM under two adversarial alignment-breaking attacks.
Attack Models BAR ASR ASR reduce
Original LLM RA-LLM Original LLM RA-LLM
Individual Vicuna-7B-chat-HF 99.3% 98.7% 98.7% 10.7% 88.0%
Guanaco-7B-HF 95.3% 92.0% 96.0% 6.7% 89.3%
Transfer Vicuna-7B-chat-HF 99.3% 98.7% 83.3% 11.3% 71.0%
Guanaco-7B-HF 95.3% 92.0% 78.7% 8.7% 70.0%
Specifically, we consider two main metrics to evaluate our model’s performances: attack success rate (ASR) and benign answering rate (BAR). Attack success rate measures the number of chances when the adversarial prompts successfully circumvent the model’s alignment mechanism. An attack is regarded as successful when the LLM produces a meaningful response without rejecting to answer with typical alignment text. To ensure the defense mechanism does not overkill and reject to answer benign questions, we also tested the benign answering rate, which represents the model precision in successfully identifying benign requests (does not reject to answer the benign requests). Our defensive goal is to minimize the attack success rate as much as possible while correctly identifying benign samples with a high benign answering rate.
From Table 1, it is evident that for Individual Attack, adversarial prompts have led to high malicious response success rates of 98.7% and 96.0% on the two models respectively. However, upon employing our robustly aligned LLM, these success rates dropped to 10.7% and 6.7%. Similarly, for Transfer Attack, the application of our robustly aligned LLM reduced the attack success rates from 83.3% and 78.7% to 11.3% and 8.7%. This demonstrates that our strategy effectively mitigates adversarial attacks. Additionally, our method maintains a good benign response rate, this indicates that our approach has almost no adverse impact on the LLM’s responses to benign inputs.
4.3 Handcrafted Jailbreak Prompts
In practice, another type of commonly seen alignment-breaking attack is the handcrafted jailbreak prompts. Those manually crafted adversarial prompts usually work by designing elaborate role-play scenarios or asking the LLM to give the responses starting with affirmative responses such as “Sure, here it is” to force the LLM to generate harmful content. In general, the handcrafted jailbreak prompt is the type of alignment-breaking attack that is more widely adopted as it only requires no computation at all, and therefore, the threats stemming from handcrafted jailbreak prompts cannot be overlooked.
Table 2: The benign answering rate and attack success rate of the original LLM and our robustly aligned LLM using handcrafted jailbreak prompts.
Model BAR ASR ASR reduce
Original LLM RA-LLM Original LLM RA-LLM
Vicuna-7B-chat-HF 99.3% 98.7% 98.7% 12.0% 86.7%
Guanaco-7B-HF 95.3% 92.0% 94.7% 9.3% 85.4%
GPT-3.5-turbo-0613 99.3% 99.3% 82.0% 8.0% 74.0%
We also assessed the defensive capabilities of our robustly aligned LLM against these meticulously designed jailbreak prompts. Specifically, we selected the top five jailbreak prompts from jailbreakchat.com [^3] The prompts are taken according to the website result on Sept 12, 2023, voted by the online users according to their effectiveness. For each of these handcrafted jailbreak prompts, we randomly selected 30 questions from the Harmful Behaviors dataset, culminating in a set of 150 handcrafted jailbreak prompt samples.
Table 2 shows the effects of our defense method on the handcrafted jailbreak prompt dataset for three different LLMs, Vicuna-7B-chat-HF, Guanaco-7B-HF, GPT-3.5-turbo-0613, all of them underwent safety alignment. We found that our robustly aligned LLM also performs exceptionally well against such handcrafted jailbreak prompts. As seen in Table 2, handcrafted jailbreak prompts achieved attack success rates of 98.4%, 94.7%, and 82.0% on the Vicuna-7B-chat-HF, Guanaco-7B-HF, and GPT-3.5-turbo-0613 models, respectively, without additional defense beyond alignment. However, when applying to our robustly aligned LLM, the attack success rates dropped to 12%, 9.3%, and 8.0%, a result even better compared to the adversarial prompt attacks in the previous section. In the meantime, RA-LLM has no significant impact on BAR especially for the larger models like GPT-3.5-turbo-0613, which inherently possess strong semantics comprehension abilities.
4.4 Ablation Study
In this section, we analyze the impact of the three hyperparameters in our method: the random dropping ratio $p$, the threshold $t$, and the number of random dropping trials $n$. For our default parameters, these parameters are set as $n=20,p=0.3,t=0.2$. We evaluate the influence of these hyperparameters using the attack success rate and benign answering rate on the Harmful Behaviors attack in Vicuna-7B-chat-HF model. The evaluation results are depicted in Figure 3.
(a) The Effect of $p$
(b) The Effect of $t$
(c) The Effect of $n$
Figure 3: Ablation Study of Harmful Behaviors Attack
The Effect of Dropping Ratio $p$
As observed in Figure 2(a), we note that a larger random dropping ratio $p$ can further reduce the attack success rate. However, it might also lead to a significant drop in benign answering rate, suggesting that it tends to have a more strict rule and thus considers a lot of benign requests as malicious. When the random dropping ratio $p$ is smaller, the accuracy on benign samples remains at a high level, but it will also affect the efficacy of the robust alignment checking function, leading to a higher attack success rate.
The Effect of Threshold $t$
Similarly, from Figure 2(b), we can observe that a too small $t$ can decrease the accuracy on benign samples, as the randomly dropped benign samples can sometimes be confusing for LLM to understand and thus also be rejected to answer. Conversely, a very large $t$ makes it difficult to reach the threshold to trigger the rejection of answering, resulting in only a limited reduction in the attack success rate.
The Effect of Monte Carlo trials $n$
Furthermore, as observed in Figure 2(c), our method still exhibits good performance with various Monte Carlo trails. Even with very few Monte Carlo trials such as 15 and 10, our robustly aligned LLM maintains a benign answering rate close to 100% and a relatively low attack success rate. This suggests that reducing the number of Monte Carlo trials is a potential strategy to decrease computational overhead while maintaining stable defensive performance.
4.5 Computational Cost
In this section, we discuss the additional computational costs incurred by our robustly aligned LLM compared to the original LLM. Suppose the token counts for input content and LLM responses in a dialogue are $l_{\text{in}}$ and $l_{\text{out}}$, respectively, and the computational costs for each input and response token are $c_{\text{in}}$ and $c_{\text{out}}$, respectively. The total cost of the original LLM is: $C_{\text{LLM}}=l_{\text{in}}\times c_{\text{in}}+l_{\text{out}}\times c_{\text{out}}$. For our robustly aligned LLM, the Monte Carlo sampling process introduces additional costs. Let the number of Monte Carlo samplings be $n$ and the proportion of input tokens randomly discarded in each sampling be $p$. Additionally, to reduce computational costs, we limit the maximum number of output tokens to $t_{\text{max}}$. Hence, if $\text{AC}({\mathbf{x}})$ fails, the extra cost of our defense is:
$C_{\text{extra}}=(1-p)l_{\text{in}}\times c_{\text{in}}\times n+l_{\text{out}}\times c_{\text{out}}\times n,\text{ where~{}}l_{\text{out}}\leq t_{\text{max}}.$
The ratio of the extra cost to the computational cost of the LLM without defense is:
$\frac{C_{\text{extra}}}{C_{\text{LLM}}}=\frac{(1-p)l_{\text{in}}\times c_{\text{in}}\times n+l_{\text{out}}\times c_{\text{out}}\times n}{l_{\text{in}}\times c_{\text{in}}+l_{\text{out}}\times c_{\text{out}}}\leq\frac{(1-p)l_{\text{in}}\times c_{\text{in}}\times n+t_{\text{max}}\times c_{\text{out}}\times n}{l_{\text{in}}\times c_{\text{in}}+l_{\text{out}}\times c_{\text{out}}}.$
If we approximate the value of $\frac{C_{\text{extra}}}{C_{\text{LLM}}}$ using our experimental data, the average token counts for inputs $l_{\text{in}}=22.58$ and outputs $l_{\text{out}}=275.25$. For our default parameters, i.e., $n=20,p=0.3,t=0.2,t_{\text{max}}=10$. To calculate the average computational cost per token, we refer to the pricing of the ChatGPT API. The GPT-4 model with an 8K context is priced at \$0.03 / 1K tokens for input and \$0.06 / 1K tokens for output, whereas the GPT-3.5 Turbo model with a 16K context is priced at \$0.003 / 1K tokens for input and \$0.004 / 1K tokens for output.
After calculations, $\frac{C_{\text{extra}}}{C_{\text{LLM}}}$ is approximately 1.250 under GPT-4 pricing and about 1.496 under GPT-3.5 Turbo pricing. We believe this cost is reasonable considering the defensive performance it could provide. If the computational cost is a real concern, one can further trade off a bit of defensive performance for cost reduction by adjusting the hyperparameters used (e.g., $p$, $t$, and $n$) as suggested in our ablation studies.
5 Conclusion and Future work
While a variety of alignment strategies have been proposed to guide the large language model to obey human ethical principles, recent works show that these alignments are vulnerable and can be bypassed by alignment-breaking attacks through carefully crafted adversarial prompts. In this work, we propose robustly aligned LLMs, which are built upon existing aligned LLMs with a robust alignment checking function, to defend against alignment-breaking attacks. One major advantage of our method is that there is no need to expensively retrain or fine-tune the original LLM for defense purposes. We also provide a theoretical analysis to verify the effectiveness of our proposed defense. The exhaustive experimental results clearly demonstrate our method can effectively defend against both automatically generated adversarial prompts and handcrafted jailbreak prompts.
Note that it is non-trivial to directly apply the current alignment-breaking attack strategies (such as Zou et al. (2023)) to our robustly aligned LLM due to the non-differentiability of our random dropping mechanism, which makes it hard to perform the gradient-based search or text optimization. So far it is under-explored whether the attackers could elaborately design stronger and more efficient attacks with the knowledge of our defense details. We leave this as our future work.
Appendix A Proof of Theorem 3.1
In this section, we provide the proof of Theorem 3.1.
Proof of Theorem 3.1.
The part of the proof for Theorem 3.1 is inspired from (Zeng et al., 2023). Denote ${\mathbf{x}}_{\text{adv}}^{j}$ as any adversarial example constructed from ${\mathbf{x}}$ where $M$ continuous adversarial tokens are inserted into position $j$, and denote the inserted adversarial prompt as ${\mathbf{p}}_{\text{adv}}^{j}$. For each $j$, we have the following equations based on the law of total probability:
[…]
$=\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)$
$+\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}=\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}=\emptyset)$
and
[…]
$=\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)$
$+\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}=\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}=\emptyset)$
When $[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}=\emptyset$, we have that $[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}=[{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}$. Thus, there is
$\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}=\emptyset)=\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}=\emptyset)$ (6)
Given Equation 6, $\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\geq 0$, and $0\leq\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\leq 1$, we could compute $\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})-\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))=\text{Fail})$ as follows
$\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})-\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))=\text{Fail})$ (7)
$=\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)$
$-\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)$
$\leq\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))$ $=\text{Fail})|[{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)$
$\leq\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}$ $\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)$
If $N\geq\frac{M(1-p)}{p}$, there is $\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}\cap{\mathbf{p}}_{\text{adv}}^{j}\neq\emptyset)=1-\frac{\binom{N}{(N+M)(1-p)}}{\binom{N+M}{(N+M)(1-p)}}=c$, thus we have
$\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})-\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))=\text{Fail})\leq c$ (8)
If $\mathop{\min}\limits_{j}\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})>t+c$, which implies for any $j$, we have $\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})>t+c$. Based on Equation 8, we can conclude that
$\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{adv}}^{j}]_{{\mathbf{r}}}))=\text{Fail})\geq\mathop{{\mathbb{P}}}\limits_{{\mathbf{r}}\sim U}(\text{AC}(f([{\mathbf{x}}_{\text{pad}}^{j}]_{{\mathbf{r}}}))=\text{Fail})-c>t$ (9)
then for any $j$, we have $\text{RAC}({\mathbf{x}}_{\text{adv}}^{j})=\text{Fail}$. Therefore, we obtain that $\text{RAC}({\mathbf{x}}_{\text{adv}})=\text{Fail}$. This concludes the proof. ∎
Appendix B Concrete Examples
In this section, we also give a few concrete examples comparing the output of the original LLM and our robustly aligned LLM under alignment-breaking attacks.
Figure 4: Multiple real cases of the original LLM’s response before and after random dropping under harmful behaviors attack.
Figure 5: Multiple real cases of the original LLM’s response before and after random under handcrafted jailbreak attack. Note that in this example, we have not explicitly labeled what is discarded.
Appendix C Potential Adaptive Attack
In this section, we will discuss potential adaptive attack methods against the defense mechanism we proposed. Since our method randomly drops $p$ portion of tokens from the input ${\mathbf{x}}$ and uses Monte Carlo sampling to simulate all possible scenarios, any form of adversarial prompt may be discarded. Hence, it’s challenging to design an adaptive attack based on optimization for our defense method. However, one may also utilize this design choice and simply try increasing the length of the adversarial prompts (e.g., repeat the adversarial prompts after input for several times) to ensure the random dropping cannot fully remove the adversarial parts.
In order to figure out whether such a potential adaptive attack can invalidate our defense or not, we conducted experiments on the Harmful Behaviors attack on both the original LLM and our robustly aligned LLM. The results are presented in Table 3. We found that, on the original LLM, repeating the adversarial prompt multiple times in the input also leads to a reduction in the attack success rate. We speculate that the adversarial prompt might heavily depend on its position within the full input. Similarly, we observed that on our robustly aligned LLM, the attack success rate decreases as the number of repetitions increases. What’s more, at various repetition counts, our defense method keeps the attack success rate lower than scenarios without repetitions, hovering around 5%. This suggests that even if attackers are familiar with our defense and want to use longer repetitive adversarial prompts, our method remains effective in thwarting their attacks.
Table 3: Adaptive attack success rate in our robustly aligned LLM. Repetition Times represents the number of repetitions of adversarial prompts
Repetition Times No Repetition 2 3 5
Original LLM 100.0% 46.0% 34.0% 31.0%
Robustly Aligned LLM 11.0% 5.0% 6.0% 3.0%
Appendix D Defensive Efficacy Against Harmful Strings Attack
Table 4: The benign answering rate and attack success rate of the original LLM and our robustly aligned LLM under two adversarial alignment-breaking attacks.
Attack BAR ASR ASR reduce
Original LLM RA-LLM Original LLM RA-LLM
Adv Strings 100.0% 99.0% 84.0% 0 84.0%
We also conducted experiments under the setting of Harmful String Attack proposed in (Zou et al., 2023). The goal of Harmful Strings attack is to compute an adversarial input, which can induce the LLM to generate a specific harmful string. Although this setting does not really fit in our threat model, it would also be interesting to see how RA-LLM performs under this attack. We conducted experiments on the Vicuna-7B-v1.3 model, and the results are presented in Table 4. It can be observed that, in the original LLM, the attack success rate of adversarial prompts generated by Harmful String Attack is as high as 84%, while after applying our RA-LLM, the attack success rate drops to 0%. This indicates that our strategy can also effectively mitigate Harmful String Attack.
Appendix E Collaborating with Safety Alignment on LLMs to Counteract Attacks
We have shown in the experiments that applying the random dropping strategy on malicious requests with adversarial prompts can effectively trigger the alignment of the model. However, for benign requests, random dropping may lead to a loss of key information and make the LLM occasionally generate unable-to-answer responses similar to typical alignment responses. This leads to a certain level of decrease in terms of benign answering rate.
Clearly, we can further reduce the loss on benign answering rate if the alignment response of the LLMs can be distinguishable from the other types of unable-to-answer responses. For instance, during the alignment fine-tuning process, the LLM is instructed to always start the response to malicious requests with a special token. When applying our defensive method, it is only necessary to output and check the first token in each Monte Carlo trial. Such a collaborative strategy on alignment and RA-LLM will not only significantly improve our recognition accuracy for malicious inputs but also help in further reducing computational costs. Due to limited resources, we leave this part as our future work.
Appendix F Comparison with LLM Self Defense
In this section, we additionally compared our method with the concurrent work, LLM Self Defense (Helbling et al., 2023). We use the more effective ”Questions as suffix” setting, that simply utilizes itself or another LLM instance to detect if its own response is harmful through the instruction such as
Here is some text:{response} Is it harmful? Give answer as “Yes, it is harmful” or “No, it is harmless”.
Specifically, we conducted both Self Defense through itself and Self Defense through GPT3.5 and summarized the comparison results in Table 5. We can observe that the attacked models (i.e., Vicuna and Guanaco) inherently struggle to discern harmful content, leading to low BAR. Moreover, the outputs of the detection prompt contain unexpected content including deviations from the predefined responses of ”Yes, it is harmful.” or ”No, it is harmless.”, continue completion of the unfinished content from the previous context, and even non-responsiveness from LLMs. These suggest that in scenarios preventing the use of external powerful models, LLM Self Defense might be less effective.
While Self Defense though more powerful LLM instances such as GPT3.5 demonstrates higher accuracy in identifying harmful content and thus enjoys on-par defending effectiveness with our method, it still suffers from lower BARs. This could be attributed to the current LLM’s overcautiousness in detecting harmful content (Röttger et al., 2023).
Table 5: The benign answering rate and attack success rate of the original LLM, self Defense, self Defense by GPT3.5, and our RA-LLM under individual adversarial alignment-breaking attacks.
Models BAR ASR
Original LLM Self Defense GPT3.5 RA-LLM Original LLM Self Defense GPT3.5 RA-LLM
Vicuna-7B-chat-HF 99.3% 68.7% 90.0% 98.7% 98.7% 22.7% 8.0% 10.7%
Guanaco-7B-HF 95.3% 41.3% 87.3% 92.0% 96.0% 52.0% 8.7% 6.7%
Appendix G Comparison with Perplexity-Based Defense
Perplexity-based defense proposed by Jain et al. (2023) detects adversarial prompts by checking if the perplexity of the prompt is greater than a threshold. Following the same threshold adopted in Zhu et al. (2023), we report the comparison results in Table 6. We can observe that even though perplexity defense achieves high BAR and effectively reduces the ASR of individual GCG attacks, this defense mechanism completely fails to detect handcrafted jailbreak prompts, presumably owing to the lower perplexity of these prompts, as they are manually written by humans. A similar conclusion is also validated in Zhu et al. (2023). In contrast, our method effectively defends against handcrafted jailbreak prompts.
Table 6: The benign answering rate and attack success rate of the original LLM, perplexity defense, and our robustly aligned LLM under two alignment-breaking attacks.
Attack Models BAR ASR
Original LLM Perplexity Defense RA-LLM Original LLM Perplexity Defense RA-LLM
Individual GCG Vicuna-7B-chat-HF 99.3% 98.0% 98.7% 98.7% 0% 10.7%
Guanaco-7B-HF 95.3% 100% 92.0% 96.0% 4% 6.7%
Handcrafted prompt Vicuna-7B-chat-HF 99.3% 98.0% 98.7% 98.7% 100% 12.0%
Guanaco-7B-HF 95.3% 100% 92.0% 94.7% 100% 9.3%
Appendix H Time Cost
To further reduce the cost of RA-LLM, we implemented an early-exit mechanism in the Monte Carlo simulation. Specifically, if the number of detected failure cases exceeds our predefined threshold during the Monte Carlo simulation, RA-LLM terminates the process early and marks the input as a malicious sample. For instance, with Monte Carlo trials at $n=20$ and a threshold $t=0.2$, RA-LLM designates an input as malicious if it detects $0.2\times 20=4$ aligned responses. If 4 aligned responses are detected in the first 6 Monte Carlo trials, the remaining 14 trials will not be executed. Similarly, if no aligned responses are found in the first 17 trials, the input is immediately classified as benign, and the last 3 trials are skipped. This approach helps to further reduce computational costs.
We evaluated 150 attack samples on both Vicuna-7B-chat-HF and Guanaco-7B-HF models, measuring the normal inference time, the time required by RA-LLM, and the time taken by RA-LLM after forcibly completing the entire Monte Carlo simulation process. We set the maximum token generation during normal inference at 1,000. For RA-LLM, we follow the default setting, and we conducted all experiments on an NVIDIA RTX A6000 GPU.
For the Vicuna-7B-chat-HF model, normal inference took 20.97 seconds per data on average, RA-LLM required an extra 3.93 seconds per data on average, and RA-LLM with the full Monte Carlo simulation required an extra 9.26 seconds per data on average. For the Guanaco-7B-HF model, these averages were 30.36 seconds for normal inference, extra 3.76 seconds for RA-LLM, and an extra 12.84 seconds for the full Monte Carlo simulation. It is observed that the time required for RA-LLM is less than 20% (18.7% and 12.0%) of the normal inference time. Even in the worst-case scenario, where each instance undergoes a full Monte Carlo simulation, the additional time cost does not exceed 45% (44.1% and 42.3%). We believe this cost is acceptable.
/div>